OpenStudy (anonymous):

Use clustering to estimate the group of numbers: 9.8 + 10.1 + 11.5 + 9.2 + 9.1 40 45 50 55

OpenStudy (anonymous):

55 its closest to it if you round

OpenStudy (anonymous):

55

OpenStudy (anonymous):

MEDAL PLZ

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

thanks

OpenStudy (acxbox22):

i think it is 50 because it is 10+10+12+9+9=20+12+18=20+30=50